4. a) Evaluate : integrate (x ^ 3 + 1/(x ^ 3)) dx


b) int 0 ^ pi 2 -4 sin xdx 3+4 cos x

1
Expert's answer
2021-12-08T12:19:03-0500

4.

a)


"\int (x^3+\dfrac{1}{x^3})dx=\dfrac{x^4}{4}-\dfrac{1}{2x^2}+C"

b)


"\displaystyle\int_{0}^{\pi/2}\dfrac{-4\sin x}{3+4\cos x}dx"

"\int \dfrac{-4\sin x}{3+4\cos x}dx"

Use "u" -substitution

"u=3+4\cos x, du=-4\sin x"


"\int \dfrac{-4\sin x}{3+4\cos x}dx=\int \dfrac{du}{u}dx=\ln(|u|)+C"

"=\ln(|3+4\cos x|)+C"

"\displaystyle\int_{0}^{\pi/2}\dfrac{-4\sin x}{3+4\cos x}dx=[\ln(|3+4\cos x|)]\begin{matrix}
 \pi/2 \\
 0
\end{matrix}"

"=\ln(|3+4\cos (\pi/2)|)-\ln(|3+4\cos (0)|)"

"=\ln 3-\ln 7=\ln (3/7)"